At January 1, 2020, Betty DeRose, Inc. reported work in process totaling $18,670.
This balance was made up of three jobs having the following costs:

                                    Job #171      Job #236      Job #412
  total cost at Jan. 1, 2020 .....   $4,390        $6,370        $7,910

During 2020, Betty DeRose reported the following transactions:

a. Direct materials used were: Job #171, $6,700; Job #236, $4,100; and
   Job #412, $3,900.
b. Direct labor incurred was: Job #171, $5,000; Job #236, $2,800; and
   Job #412, $3,000.
c. Overhead was applied to jobs at a rate of 85% of direct materials used.
d. Job #171, which consisted of 100 units, was completed and transferred
   to the finished goods warehouse.
e. Job #412, which consisted of 250 units, was completed and transferred
   to the finished goods warehouse.
f. 40 units from Job #171 were sold and shipped to the customer; the 40
   units were sold at a total selling price of $12,000.
g. The actual overhead cost incurred during 2020 amounted to $10,000.

Calculate the amount of gross profit reported in Betty DeRose's income
statement for 2020.
Answer #1

Job Cost sheet For Job #171

Beginning balance 4390
Direct material 6700
Direct labor 5000
Overhead applied (6700*85%) 5695
Total Job Cost 21785
Unit 100
Cost per unit 217.85

Unadjusted cost of goods sold = 40*217.85 = 8714

Applied overhead = (6700+4100+3900)*85% = 12495

Actual overhead = 10000

Over applied overhead = 12495-10000 = 2495

Adjusted cost of goods sold = 8714-2495 = 6219

Gross profit = Sales-Cost of goods sold = 12000-6219 = 5781